OpenStudy (anonymous):

f(x)=1/4x-5 find (f(times)f^-1)(-2)

sam (.sam.):

answer 1/17

sam (.sam.):

\[ff ^{-1}(2)\] Find inverse of f(x) first, y=1/(4x-5) x=(1+5y)/y \[f ^{-1}(x)=(1+5x)/x\] Substitute into equation. =f((1+5(2))/2) =1/(4(11/2)-5) =1/17

OpenStudy (anonymous):

x All real numbers -2 0 these are the answer choices

OpenStudy (ash2326):

I get it Let me write

OpenStudy (ash2326):

We have \[f (f^{-1}(x))\] we have to find value of this for x=-2 we know that \[f f^{-1} (x)= x\] so \[f f^{-1} (-2)= -2\]

OpenStudy (anonymous):

Thank you so so sooooo much D: I've been trying to get an answer for this problem forever <3333